OpenStudy (anonymous):

Find constants a and b so that the minimum for the parabola f(x)=x^2+ax+b is at the point (3,7)

OpenStudy (anonymous):

2x+a=0 2x=-a 2*3=-a -6=a b=put x and a and y =16

OpenStudy (anonymous):

i hope this is the answer cheers

OpenStudy (anonymous):

Kropot are u actually typing?

OpenStudy (kropot72):

Differentiating the function with respect to x gives (2x + a) For a maximum or a minimum put 2x + a = 0. Solving for a gives a = -6. Further substitution gives b = 16